OpenStudy (anonymous):

Which complex number has an absolute value of 5? A. –3 + 4i B. 2 + 3i C. 7 – 2i D. 9 + 4i

OpenStudy (anonymous):

d

OpenStudy (tkhunny):

"Absolute Value" is sort of unusual language for the modulus of a Complex Number. In any case, the Pythagorean Theorem solves this for you. D. 9 + 4i ==> 9^2 + 4^2 = 81 + 16 = 97 ==> |9 + 4i| = \(\sqrt{97} > 5\) C. It should be obvious that one leg of your right triangle greater than 5 is going to have a modulus greater than 5. B. 2 + 3i ==> 2^2 + 3^2 = 4 + 9 = 13 ==> |2 + 3i| = \(\sqrt{13} < 5\) A. Let's see what you get on this one.

OpenStudy (anonymous):

Class A Class B Class C 80 72 100 85 90 90 87 85 94 92 84 100 86 91 75 90 78 30 94 94 94 100 86 72 75 82 85 45 78 87 85 86 42 38 78 100 72 95 90 94 90 94 62 88 60 Question 1 Analyze the data for Class A and Class B. Part 1: What are the means and standard deviations for Classes A and B? Part 2: Make a box plot of the data for Classes A and B. Identify the median, first and third quartiles, and minimum and maximum data points. Part 3: Explain how the values for Classes A and B vary about the mean and median. What information does this give Mrs. Smith? Question 2 Analyze and compare the data for Class B and Class C. Part 1: Make a dot plot for each of Classes B and C. Part 2: What information is easily seen from the dot plots for Classes B and C? Part 3: Look at the box plot and dot plot created for Class B. What are the strengths and weaknesses of each type of plot? Part 4: Which of the two classes (B or C) performed better on the test? Justify your conclusion. Question 3 Mrs. Smith gives three quizzes before each test. She allows each student to drop their lowest test grade during the grading period and replace it with the average of their quiz grades. The grades in red below were replaced with the average quiz grades. Class A Class B Class C 80 90 100 85 90 90 87 85 94 92 84 100 86 91 75 90 78 84 94 94 94 100 86 72 75 82 85 70 78 87 85 86 72 38 78 100 72 95 90 94 90 94 82 88 60 Part 1: How does this change the data for Classes A and C? Part 2: Based upon these changes, which of the three classes did the best on the exam? Justify your answer. SOME ONE HELP